OpenStudy (anonymous):

A ribbon of nichrome (resistivity = 1.00 × 10-6 Ω·m) with a rectangular cross sectional area of (0.25 mm × 1.00 mm) will be used as a heating element in a toaster. How long should this ribbon be, if the total resistance is intended to be 1.5 Ω? 1.24 cm 2.79 cm 37.5 cm 12.4 cm 21.9 cm

OpenStudy (kropot72):

The cross-sectional area of the ribbon in square meters is given by \[c.s.a=0.25\times 10^{-3}\times 1.00\times 10^{-3}\] The required resistance is given by \[resistance=\frac{resistivity \times length}{c.s.a.}\] \[1.5=\frac{1.00\times 10^{-6}\times l}{0.25\times 10^{-3}\times 1.00\times 10^{-3}}\] from which we get \[l=1.5\times 0.25\times 100=?\ cm\]